notmeta:

does this equation have a solution 6 – 1.5(4b + 8) = 3(-6 – 2b)

notmeta:

I think it is but I'm not sure and don't wanna get it wrong

Breathless:

toga got this-

notmeta:

I also have another one easier tho "Michelle correctly solved the equation 4x - 6 = 4(x – 2) + 1 for the variable x, and the last step of her work was: 0 = 1. Select (by clicking it) the statement that best describes the solution to the equation Michelle was solving?"

toga:

Let's solve it step by step: 6 - 1.5(4b + 8) = 3(-6 - 2b) First, let's distribute the 1.5 and 3: 6 - 6b - 12 = -18 - 6b Combine like terms: -6b - 6 = -18 - 6b Now, adding 6b to both sides: -6 = -18 This equation leads to a contradiction, -6 is not equal to -18. So, the equation has no solution.

notmeta:

thanks toga

toga:

@notmeta wrote:
I also have another one easier tho "Michelle correctly solved the equation 4x - 6 = 4(x – 2) + 1 for the variable x, and the last step of her work was: 0 = 1. Select (by clicking it) the statement that best describes the solution to the equation Michelle was solving?"
here is how to solve it: First, distribute the 4 on the right side of the equation: 4x - 6 = 4x - 8 + 1 Then, combine like terms: 4x - 6 = 4x - 7 Now, subtract 4x from both sides: -6 = -7 This leads to a contradiction, showing that there is no solution to this equation. Therefore, the correct statement that best describes the solution to the equation Michelle was solving is "The equation has no solution."

Breathless:

see? Toga so smart

notmeta:

literally my hero

toga:

@breathless wrote:
see? Toga so smart
everybody is smart in their own ways mine happens to be knowing how to solve math problems and remembering useless information about history, English, and other things

Breathless:

@toga wrote:
@breathless wrote:
see? Toga so smart
everybody is smart in their own ways mine happens to be knowing how to solve math problems and remembering useless information about history, English, and other things
lol

Breathless:

anyways I gotta finish my bath Brb